OpenStudy (anonymous):

The ratio of the distances a 7-iron and a 5-iron will drive a golf ball is 5 to 6. If a golfer averages 155 yards with a 7 iron how far should he average with a 5 iron? Question 6 options: 144 yd 186 yd 166 yd 129 yd

OpenStudy (dumbcow):

set up proportion \[\frac{5}{6} = \frac{155}{N}\] solve for N

OpenStudy (dumbcow):

\[N = 155*\frac{6}{5}\]

OpenStudy (texaschic101):

There is nothing else to say....dumbcow summed it up very well